OpenStudy (anonymous):

I NEED HELP!!!!!!!!! What is the volume of a pyramid with base area B = 9 ft2 and height h = 11 ft? A. 33 ft3 B. 49.5 ft3 C. 66 ft3 D. 99 ft3

OpenStudy (anonymous):

 -> 3*3 = Area of 9. So each side of the base is 3, and that means the width is 3. We know the height is 11. Also, we know that the formula for finding the volume of a pyramid is: V=(lwh)/3 Let's just substitute in our known values: V=((3)(3)(11))/3 Solve! V = 99/3 = 33 ft^3

OpenStudy (anonymous):

That would be A. ^_^